A flashlight operates on 9.00 V battery and has a resistance of 45.0 . At what rate does the flashlight use electric energy?
1
Expert's answer
2019-05-27T11:14:40-0400

The speed at which the flashlight uses electric energy is called the electrical power. Electrical power P can be calculated by


"P=\frac{{{V}^{2}}}{R}"

where V=9.00 volts is the voltage on battery, R=45 ohms is the resistance of flashlight. Substituting these values ​​into the ratio, we get


"P=\frac{{{\left( 9.00\,V \right)}^{2}}}{45\,\Omega }=\frac{81\,{{V}^{2}}}{45\,\Omega \,}=1.8\,\frac{J}{s}=1.8\,W"

So the speed at which the flashlight use the electrical energy is 1.8 joules per second, or 1.8 watts